# The FinBen: An Holistic Financial Benchmark for Large Language Models

### v1

@@ 4.2 Advanced Cognitive Engagement Tasks Analysis
In the text generation task, Gemini emerges as the frontrunner on the EDTSUM abstractive text summarization dataset, illustrating its prowess in generating coherent summaries. Nevertheless, all models face challenges with extractive summarization, which demands the generation of precise label sequences for sentences. In the forecasting task, Gemini distinguishes itself across most datasets, except in the Australian credit scoring dataset, where GPT-4 demonstrates superior performance.
Among open-source LLMs, LLaMA2 70B stands out in text summarization, whereas LLaMA2-7B-chat excels in forecasting tasks. Despite instruction tuning with datasets like BigData22 and ACL18, FinMA 7B lags behind peers such as Falcon 7B in forecasting performance, underscoring the need for more effective improvement strategies. CFGPT sft-7B-Full consistently shows a decrease in performance compared to its foundational model, InternLM 7B. For forecasting, it is crucial to acknowledge that all LLMs do not meet expected outcomes and fall behind traditional methodologies. This consistent observation with existing studies Feng et al. (2023); Xie et al. (2023b) underlines a notable deficiency in LLMs’ capacity to tackle advanced cognitive tasks as effectively as conventional methods.
This analysis reveals significant potential for enhancement in LLMs, including industry leaders like GPT-4 and Gemini, particularly in text generation and forecasting tasks that demand higher cognitive skills.
4.3 General Intelligence Tasks Analysis
The comparative analysis of various Large Language Models (LLMs) on the complex task of stock trading, which demands a high degree of general intelligence, is presented in Table 4555For detail trading performance, please see Appendix E. The results indicate a superior performance of all LLMs over the traditional Buy & Hold strategy, highlighting their efficacy in formulating more advantageous trading decisions. Among the evaluated LLMs, GPT-4 distinguishes itself by attaining the highest Sharpe Ratio (SR), exceeding 1. This achievement underscores GPT-4’s proficiency in optimizing profit against the risk, a capability that appears somewhat diminished in other LLMs, which tend to expose investors to higher risk for lesser returns. Additionally, GPT-4 demonstrates the minimal Max Drawdown (MDD), suggesting that it limits potential losses more effectively than its counterparts, thereby offering a more secure investment avenue.
@@ 4.3 General Intelligence Tasks Analysis
In contrast, ChatGPT exhibits significantly lower performance metrics, indicating limitations in its financial decision-making capabilities. Gemini, on the other hand, secures the position of second-best performer, showcasing lower risk and volatility in comparison to GPT-4, yet maintaining commendable returns. When considering open-source models, it is observed that LLaMA-70B, despite its lower volatility, yields the least profit among the LLMs, highlighting a trade-off between risk management and profitability.
@@
For smaller models with parameters less than 70 billion, a marked inability to adhere to trading instructions consistently across transactions is noted, attributed to their limited comprehension, extraction capabilities, and constrained context windows. This limitation underscores the critical challenges smaller LLMs face in tasks requiring intricate financial reasoning and decision-making, thereby spotlighting the necessity for more advanced models to tackle such high-level cognitive tasks effectively.
In essence, the exceptional performance of LLMs in the stock trading task illuminates their capacity to embody general intelligence within the financial domain. This capacity, rooted in the integration of diverse cognitive skills and the application of these skills to real-world financial challenges, heralds a new era of financial analysis and decision-making. Our findings, thereby, not only affirm the significant potential of LLMs in navigating the complexities of financial markets but also suggest a promising trajectory for their further development and application in tasks demanding a high level of general intelligence.
@@ 5 Conclusion
In this work, we introduce a comprehensive financial benchmark the FinBen specifically designed for evaluating LLMs in the financial domain. This benchmark encompasses 35 diverse datasets from 23 tasks, organized into three spectrums of difficulty. Unlike previous benchmarks in the financial domain, the FinBen extends its evaluation to encompass a broad spectrum of tasks, including quantification, extraction, understanding, generation, forecasting. Notably, for the first time, it incorporates a direct trading task through an agent-based evaluation framework. Our comprehensive evaluation of 15 representative LLMs yields several key insights: 1) GPT-4 emerges as the top performer in tasks related to quantification, extraction, understanding, and trading, whereas Gemini leads in generation and forecasting tasks. 2) While existing LLMs demonstrate commendable performance on foundational tasks, their effectiveness on more cognitively demanding tasks and those requiring general intelligence appears constrained. 3) The findings highlight the capacity of LLMs to directly inform trading decisions, suggesting a promising avenue for future research. Moving forward, we aim to expand FinBen to encompass additional languages and a wider array of financial trading tasks, further broadening the benchmark’s applicability and utility in advancing the field of financial LLMs.
Limitations
Despite the groundbreaking efforts to benchmark LLMs in the financial domain through the FinBen, we acknowledge several inherent limitations that could impact the benchmark’s effectiveness and applicability:
Dataset Size Limitations: A primary challenge faced in the development of the FinBen is the restricted size of available datasets, a common issue in the niche field of open-source financial data. This limitation may affect the depth of the models’ financial understanding and their ability to generalize across the full spectrum of financial contexts. Model Size Limitations: Due to computational resource constraints, our evaluation was limited to the LLaMA 70B model. This restriction potentially overlooks the capabilities and performance nuances that larger or differently architected models might demonstrate on FinBen’s comprehensive task suite. Generalizability: The tasks, particularly those involving trading and forecasting, are predominantly based on data from American markets and English-language texts. This focus may limit the benchmark’s applicability to global financial markets, where linguistic diversity and unique market dynamics play a crucial role. Potential Negative Impacts: While the FinBen aims to propel the field of financial language understanding forward, it is crucial to consider the potential for misuse, such as the propagation of financial misinformation or the exertion of unethical influence on markets. These risks underscore the importance of responsible usage and further safeguards in the deployment of LLMs trained or evaluated with the FinBen [^6] For a detailed ethical and legal statement concerning this work, please see Appendix..
Ethical Statement
The development and dissemination of the FinBen by the authors carry full responsibility for any potential violation of rights or arising legal issues. Diligent efforts have been undertaken to ensure the construction of the FinBen respects privacy and conforms to established ethical guidelines. The datasets compiled within FinBen are shared under the MIT license, with the expectation that users agree to adhere to its conditions.
@@ Appendix A Contributions
Science Leadership: Qianqian Xie, Min Peng, Sophia Ananiadou, Alejandro Lopez-Lira, Hao Wang, Yanzhao Lai, Benyou Wang, Xiao-yang Liu, Gang Hu, Jiajia Huang, Jimin Huang. Contributors: Weiguang Han, Zhengyu Chen, Ruoyu Xiang, Xiao Zhang, Yueru He, Mengxi Xiao, Dong Li, Yongfu Dai, Duanyu Feng, Yijing Xu, Haoqiang Kang, Ziyan Kuang, Chenhan Yuan, Kailai Yang, Zheheng Luo, Tianlin Zhang, Zhiwei Liu, Guojun Xiong, Zhiyang Deng, Yuechen Jiang, Zhiyuan Yao, Haohang Li, Yangyang Yu
@@ Appendix B Other LLMs Performance
Dataset Metrics Baichuan 7B CodeLLaMA 7B DISC- FinLLM
FPB F1 0.36 0.34 0.29
Acc 0.32 0.39 0.26
FiQA-SA F1 0.17 0.66 0.32
@@
Headlines AvgF1 0.60 0.60 0.60
FOMC F1 0.16 0.14 0.19
Acc 0.25 0.27 0.28
FinArg-ACC MicroF1 0.34 0.28 0.29
FinArg-ARC MicroF1 0.17 0.25 0.29
MultiFin MicroF1 0.06 0.21 0.29
M&A MicroF1 0.02 0.54 0.29
MLESG MicroF1 0.00 0.10 0.29
NER EntityF1 0.00 0.07 0.12
FINER-ORD EntityF1 0.00 0.00 0.00
FinRED F1 0.00 0.00 0.00
SC F1 0.57 0.85 0.00
CD F1 0.00 0.00 0.00
FinQA EmAcc 0.00 0.00 0.00
TATQA EmAcc 0.00 0.00 0.00
ConvFinQA EmAcc 0.00 0.00 0.00
FNXL EntityF1 0.00 0.00 0.00
FSRL EntityF1 0.00 0.00 0.00
EDTSUM Rouge-1 0.22 0.10 0.22
BertScore 0.54 0.67 0.61
BartScore -4.57 -3.62 -4.13
ECTSUM Rouge-1 0.00 0.00 0.00
BertScore 0.00 0.00 0.00
BartScore -5.18 -5.18 -5.18
BigData22 Acc 0.53 0.52 0.44
MCC -0.01 -0.01 -0.05
ACL18 Acc 0.50 0.51 0.50
MCC 0.00 0.00 0.02
CIKM18 Acc 0.53 0.51 0.44
MCC -0.05 0.02 -0.03
German F1 0.52 0.66 0.52
MCC 0.00 0.00 0.00
Australian F1 0.26 0.43 0.26
MCC 0.00 0.00 0.00
LendingClub F1 0.72 0.81 0.72
MCC -0.01 0.00 0.00
ccf F1 0.97 0.00 0.66
MCC 0.00 0.00 -0.04
ccfraud F1 0.00 0.06 0.46
MCC 0.00 0.00 0.02
polish F1 0.91 0.47 0.92
MCC 0.02 0.04 0.00
taiwan F1 0.70 0.36 0.95
MCC -0.02 -0.03 0.00
portoseguro F1 0.01 0.88 0.63
MCC 0.01 -0.01 -0.02
travelinsurance F1 0.03 0.02 0.00
MCC -0.09 0.00 0.00
Table 5: The zero-shot and few-shot performance of other LLMs on the FinBen.
@@ Appendix C Instructions
Table 6: The prompt of each dataset.
@@
Table 7: The example prompt for each dataset. FiQA-SA has two types of text, including news headlines and tweets. We will fill the detailed text type into {category} for each data sample. For stock movement prediction data such as BigData22, we will fill {tid} and {point} with the detailed stock name and time from each data sample.
@@ D.1 Financial Large Language Models
Recent years have seen a significant surge in research on finance-specific LLMs, expanding on the groundwork laid by general-purpose language models Lee et al. (2024); Liu et al. (2023b); Xie et al. (2023a); Zhang et al. (2024); Dai et al. (2024). Financial pre-trained language models (FinPLMs) like FinBERT Araci (2019); Yang et al. (2020b); Liu et al. (2020), derived from BERT, and FLANG Shah et al. (2022), based on ELECTRA, have been developed using domain-specific data for enhanced performance in tasks like sentiment analysis and stock prediction. The open-source release of Meta AI’s LLaMA Touvron et al. (2023) has fueled further innovation in Financial LLMs (FinLLMs), with models like FinMA Xie et al. (2023b), InvestLM Yang et al. (2023b), and FinGPT Wang et al. (2023); Liu et al. (2023a) leveraging advanced tuning strategies Zhang et al. (2023a) for financial applications. BloombergGPT Wu et al. (2023) stands out as a BLOOM-based, closed-source models tailored for the financial industry. Additionally, the Chinese financial sector has seen the emergence of models like XuanYuan 2.0 Zhang et al. (2023c), integrating broad and specialized knowledge, FinBART Hongyuan et al. (2023) for financial communication, and CFGPT Li et al. (2023a), which includes a comprehensive dataset for targeted pre-training and fine-tuning.

### v2

@@ 4.2 Question Answering & Text Generation Results
In the TG tasks, Gemini emerges as the frontrunner on the EDTSUM abstractive text summarization dataset, illustrating its prowess in generating coherent summaries. Nevertheless, all models face challenges with extractive summarization, which demands the generation of precise label sequences for sentences. Among open-source LLMs, LLaMA2 70B stands out in text summarization. Conversely, CFGPT sft-7B-Full consistently shows a decrease in performance compared to its foundational model, InternLM 7B.
4.3 Forecasting & Risk Management Results
For forecasting, it is crucial to acknowledge that all LLMs fail to meet expected outcomes and lag behind traditional methodologies. This consistent observation with existing studies Xie et al. (2023b) underlines a notable deficiency in LLMs’ capacity to tackle forecasting as effectively as traditional methods. Even the best-performing models, such as GPT-4 and Gemini, only perform slightly better than random guessing. This reveals significant potential for enhancement in LLMs, including industry leaders like GPT-4 and Gemini, particularly in forecasting tasks that demand complex reasoning abilities.
In RM tasks, such as credit scoring, fraud detection, and identifying financial distress, data often exhibit significant imbalances. Instances representing individuals with low credit scores, those prone to fraud, and companies at risk of financial distress constitute only a small percentage of the overall dataset. In such scenarios, LLMs with low instruction-following abilities (such as LLaMA2-7B-chat and LLaMA2-70B) tend to classify all cases into a single class, resulting in an MCC score of 0. These tasks, with tabular inputs and highly imbalanced distribution, pose a significant challenge for LLMs in the financial domain.
4.4 Decision Making Results
The comparative analysis of various LLMs on the complex task of stock trading, is presented in Table 4666For detailed trading performance, please see Appendix E. This task requires models to understand, summarize, and reason with multimodal financial data (texts and time series), leading to sophisticated trading decisions that necessitate a range of skills, from fundamental comprehension and summarization to reasoning and decision-making.
Among the evaluated LLMs, GPT-4 distinguishes itself by achieving the highest Sharpe Ratio (SR) over 1, indicating superior investment performance through optimal risk-return balance. It also records the minimal Maximum Drawdown (MDD), suggesting effective limitation of potential losses, thereby offering a more secure investment avenue compared to other models, including those using reinforcement learning methods like DQN, PPO, and A2C, which show significantly lower SR and higher MDD.
Tables 4 and 9 reinforce these findings, highlighting GPT-4’s exceptional performance in this challenging domain. Additional results and analyses from these models in Table 5 contrast their performances with the traditional Buy & Hold strategy, which considerably lags behind.
Table 4: The average trading performance (95% Confidence Interval) comparison for different LLMs across 10 stocks. The results include large LLMs only ($\geq 70B$), as models with smaller contexts have difficulty understanding the instructions and producing a static strategy of holding.
[…]
Buy & Hold -4.00 $\pm$ 22.39 0.02 $\pm$ 0.87 3.59 $\pm$ 1.34 56.43 $\pm$ 21.00 30.67 $\pm$ 17.48
GPT-4 28.19 $\pm$ 25.27 1.51 $\pm$ 1.08 2.52 $\pm$ 1.30 39.88 $\pm$ 20.66 18.34 $\pm$ 9.77
GPT-4o -5.54 $\pm$ 19.12 -0.19 $\pm$ 0.84 2.73 $\pm$ 1.30 43.62 $\pm$ 20.67 29.96 $\pm$ 18.89
GPT3.5-Turbo 4.48 $\pm$ 22.23 0.15 $\pm$ 0.82 2.84 $\pm$ 1.47 45.39 $\pm$ 23.35 28.83 $\pm$ 15.40
llama2-70B 4.02 $\pm$ 24.65 0.52 $\pm$ 1.48 2.18 $\pm$ 1.28 34.86 $\pm$ 20.38 25.55 $\pm$ 16.83
llama3-70B -2.57 $\pm$ 22.63 -0.04 $\pm$ 1.19 2.71 $\pm$ 1.54 43.42 $\pm$ 24.65 29.31 $\pm$ 15.57
gemini 14.95 $\pm$ 28.04 1.03 $\pm$ 1.24 2.17 $\pm$ 1.39 34.67 $\pm$ 22.23 20.13 $\pm$ 11.36
Table 5: Traditional model performances on stock trading.
Model Cumulative Return Sharpe Ratio Standard Deviation Annualized Volatility Max Drawdown
A2C -4.2232 -0.2586 2.7522 43.6898 30.5819
PPO -0.5586 0.0085 2.7531 43.7048 28.9496
DQN -2.9924 -0.1656 2.7486 43.6319 31.78
@@ 4.4 Decision Making Results
In contrast, ChatGPT exhibits significantly lower performance metrics, indicating limitations in its financial decision-making capabilities. Gemini, on the other hand, secures the position of second-best performer, showcasing lower risk and volatility compared to GPT-4, yet maintaining commendable returns. When considering open-source models, LLaMA-70B, despite its lower volatility, yields the least profit among the LLMs, highlighting a trade-off between risk management and profitability.
@@
For smaller models with parameters less than 70 billion, a marked inability to adhere to trading instructions consistently across transactions is noted. This is attributed to their limited comprehension, extraction capabilities, and constrained context windows. This limitation underscores the critical challenges smaller LLMs face in tasks requiring intricate financial reasoning and decision-making, thereby spotlighting the necessity for more advanced models to tackle decision making tasks effectively.
@@ 5 Conclusion
In this work, we present FinBen, a comprehensive benchmark specifically designed to evaluate LLMs in the financial domain. FinBen includes 36 diverse datasets spanning 24 tasks, meticulously organized to assess LLMs across seven critical aspects: information extraction, textual analysis, question answering, text generation, risk management, forecasting, and decision-making. This breadth of coverage sets FinBen apart from existing financial benchmarks, enabling a more robust and nuanced evaluation of LLM capabilities. Our evaluation of 15 LLMs, including GPT-4, ChatGPT, and Gemini, reveals their key advantages and limitations, highlighting directions for future work. Looking ahead, FinBen aims to continuously evolve, incorporating additional languages and expanding the range of financial tasks to further enhance its applicability and impact.
Limitations: We acknowledge several limitations that could impact FinBen’s effectiveness and applicability. The restricted size of available datasets may affect the models’ financial understanding and generalization across various contexts. Computational constraints limited our evaluation to the LLaMA 70B model, potentially overlooking the capabilities of larger models. Additionally, the tasks are based on American market data and English texts, which may limit the benchmark’s applicability to global financial markets. Responsible usage and safeguards are essential to prevent potential misuse, such as financial misinformation or unethical market influence [^7] For a detailed limitation concerning this work, please see Appendix..
Ethical Statement: The authors take full responsibility for any potential legal issues arising from FinBen’s development and dissemination. All data used are publicly available, non-personal, and shared under the MIT license, adhering to privacy and ethical guidelines. This manuscript and associated materials are for academic and educational use only and do not provide financial, legal, or investment advice. The authors disclaim any liability for losses or damages from using the material, and users agree to seek professional consultation and indemnify the authors against any claims arising from its use [^8] For a detailed ethical and legal statement concerning this work, please see Appendix..
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes]
    2. (b)
      Did you describe the limitations of your work? [Yes] See Limitation (Section Limitations).
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [Yes] See Ethical Statement (Section Ethical Statement).
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes] See Ethical Statement (Section Ethical Statement).
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A]
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A]
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] See Introduction (Section 1).
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [N/A] Our benchmark only includes the evaluation process.
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [Yes] See Table 4.
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] See Experimental Settings (Section 3).
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes]
    2. (b)
      Did you mention the license of the assets? [Yes]
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [Yes] Our Introduction (Section 1) contains a link for all data used in FinBen.
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [Yes] Our dataset statistics (Table 2) contains licenses for all used datasets.
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [Yes] See Ethical Statement (Section Ethical Statement).
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A]
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A]
@@ Appendix A Contributions
Science Leadership: Qianqian Xie, Min Peng, Sophia Ananiadou, Alejandro Lopez-Lira, Hao Wang, Yanzhao Lai, Benyou Wang, Xiao-yang Liu, Gang Hu, Jiajia Huang, Jimin Huang.
Contributors: Mengxi Xiao, Dong Li, Weiguang Han, Zhengyu Chen, Ruoyu Xiang, Xiao Zhang, Yueru He, Yongfu Dai, Duanyu Feng, Yijing Xu, Haoqiang Kang, Ziyan Kuang, Chenhan Yuan, Kailai Yang, Zheheng Luo, Tianlin Zhang, Zhiwei Liu, Guojun Xiong, Zhiyang Deng, Yuechen Jiang, Zhiyuan Yao, Haohang Li, Yangyang Yu
@@ Appendix B Other LLMs Performance
Table 6: The zero-shot and few-shots performance of other LLMs on the FinBen.
Dataset Metrics Baichuan 7B CodeLLaMA 7B DISC- FinLLM ChatGLM3 6B
NER EntityF1 0.00 0.07 0.12 0.25
FINER-ORD EntityF1 0.00 0.00 0.00 0.02
FinRED F1 0.00 0.00 0.00 0.00
SC F1 0.74 0.85 0.00 0.81
CD F1 0.00 0.00 0.00 0.00
FNXL EntityF1 0.00 0.00 0.00 0.00
FSRL EntityF1 0.00 0.00 0.00 0.00
FPB F1 0.17 0.34 0.29 0.74
Acc 0.23 0.39 0.26 0.74
FiQA-SA F1 0.32 0.66 0.32 0.56
@@
Headlines AvgF1 0.60 0.60 0.60 0.66
FOMC F1 0.17 0.14 0.19 0.47
Acc 0.25 0.27 0.28 0.46
FinArg-ACC MicroF1 0.36 0.28 0.29 0.25
FinArg-ARC MicroF1 0.27 0.25 0.29 0.50
MultiFin MicroF1 0.12 0.21 0.29 0.47
M&A MicroF1 0.33 0.54 0.29 0.79
MLESG MicroF1 0.04 0.10 0.29 0.16
FinQA EmAcc 0.00 0.00 0.00 0.00
TATQA EmAcc 0.00 0.00 0.00 0.07
Regulations Rouge-1 0.13 - - 0.26
BertScore 0.60 - - 0.65
ConvFinQA EmAcc 0.00 0.00 0.00 0.00
EDTSUM Rouge-1 0.02 0.10 0.22 0.13
BertScore 0.47 0.67 0.61 0.47
BartScore -6.18 -3.62 -4.13 -4.78
ECTSUM Rouge-1 0.00 0.00 0.00 0.00
BertScore 0.00 0.00 0.00 0.00
BartScore -5.18 -5.18 -5.18 -5.18
BigData22 Acc 0.53 0.52 0.44 0.47
MCC -0.01 -0.01 -0.05 0.00
ACL18 Acc 0.50 0.51 0.50 0.50
MCC -0.01 0.00 0.02 0.02
CIKM18 Acc 0.48 0.51 0.44 0.42
MCC 0.02 0.02 -0.03 0.02
German F1 0.52 0.66 0.52 0.41
MCC 0.00 0.00 0.00 -0.30
Australian F1 0.26 0.43 0.26 0.27
MCC 0.00 0.00 0.00 -0.02
LendingClub F1 0.59 0.81 0.72 0.72
MCC 0.02 0.00 0.00 -0.03
ccf F1 0.01 0.00 0.66 1.00
MCC 0.00 0.00 -0.04 0.00
ccfraud F1 0.05 0.06 0.46 0.67
MCC -0.04 0.00 0.02 -0.15
polish F1 0.92 0.47 0.92 0.05
MCC -0.01 0.04 0.00 0.00
taiwan F1 0.95 0.36 0.95 0.05
MCC 0.00 -0.03 0.00 0.01
portoseguro F1 0.00 0.88 0.63 0.95
MCC 0.00 -0.01 -0.02 0.06
travelinsurance F1 0.00 0.02 0.00 0.97
MCC 0.00 0.00 0.00 0.03
@@ Appendix C Instructions
Table 7: Quantification task datasets prompt overview.
@@
Table 8: The example prompts of remaining tasks. FiQA-SA has two types of text, including news headlines and tweets. We will fill the detailed text type into {category} for each data sample. For stock movement prediction data such as BigData22, we will fill {tid} and {point} with the detailed stock name and time from each data sample.
@@ D.1 Financial Large Language Models
Recent years have seen a significant surge in research on finance-specific LLMs, expanding on the groundwork laid by general-purpose language models (Lee et al., 2024; Liu et al., 2023b; Xie et al., 2023a; Zhang et al., 2024; Dai et al., 2024). Financial pre-trained language models (FinPLMs) like FinBERT (Araci, 2019; Yang et al., 2020b; Liu et al., 2020), derived from BERT, and FLANG (Shah et al., 2022), based on ELECTRA, have been developed using domain-specific data for enhanced performance in tasks like sentiment analysis and stock prediction. The open-source release of Meta AI’s LLaMA (Touvron et al., 2023a, b) has fueled further innovation in Financial LLMs (FinLLMs), with models like FinMA (Xie et al., 2023b), InvestLM (Yang et al., 2023b), and FinGPT (Wang et al., 2023; Liu et al., 2023a) leveraging advanced tuning strategies (Zhang et al., 2023a) for financial applications. BloombergGPT (Wu et al., 2023) stands out as a BLOOM-based, closed-source model tailored for the financial industry. Additionally, the Chinese financial sector has seen the emergence of models like XuanYuan 2.0 (Zhang et al., 2023c), integrating broad and specialized knowledge, FinBART (Hongyuan et al., 2023) for financial communication, and CFGPT (Li et al., 2023a), which includes a comprehensive dataset for targeted pre-training and fine-tuning.